OpenStudy (anonymous):

confirm that f and g are inverses by showing the f(g(x))=x and g(f(x))=x f(x)=8/x and g(x)=8/x

OpenStudy (anonymous):

f(x) = 8/x f(g(x)) = 8/8/x = 8x/8 = x g(x) = 8/x g(f(x)) = 8/8/x = 8x/8 = x

OpenStudy (anonymous):

Hence f(g(x)) = g(f(x)) they are said to be inverse